The perimeter of a standard tennis court used for doubles is 228 ft. The width is 42 ft less than the length. Find the dimensions.
Answer by ikleyn(52759)

.
L + W + L + W = 228   feet   (the perimeter)

2L + 2W = 228


L + W = 228/2 = 114.


Substitute W = L - 42.


L + (L - 42) = 114 


2L = 114 + 42 = 156


L = 156/2 = 78.


Answer.  The dimensions are  78 ft (length)  and  78 - 42 = 36 ft (width).

Solved.